Please Show Your Calculations!

Calculate the Cost of Recruitment…..

Advertising Costs = $57,000

Agency Fees = $21,000

Referral Bonuses = $3,000

Total Hires = 215

Calculate the Yield Ratio:

Resumes obtained from employment agency = 57

On-Site Interviews Granted = 15

Calculate the Acceptance Ratio:

From above example 4 candidates accepted of 9 who were offered jobs.

Calculate the Intern Conversion Ratio:

162 Interns in our program last summer

51 Interns in that supper program accepted full time positions
Answer #1

The Cost of Recruitment=(Total internal + external cost)/total number of hires

Total internal and external cost = Advertising Costs+ Agency Fees+ Referral Bonuses

Total internal and external cost = $57,000+ $21,000+ $3,000

Total internal and external cost =$81,000

Total Hires = 215

The Cost of Recruitment=($81,000)/215

The Cost of Recruitment=$376.744/Hire

Answer:- Yield Ratio = (Number of candidates invited to interview/(Number of leads from a recruitment method ) *100R

Number of leads from recruitment method=57

/ Number of candidates invited to interview=15

Yield Ratio =(15/57)*100%

Yield Ratio =26.32%

Acceptance Ratio=Candidates accepted the offer/Total candidates offered the job

Acceptance Ratio=(4/9)*100%

Acceptance Ratio=44.44%

Intern Conversion Ratio= (Total number of interns became full time employee/ Total number of interns)*100%

Intern Conversion Ratio=51/162)*100%

Intern Conversion Ratio=31.48%